OpenStudy (anonymous):

Solve 3x – 5y = 15 for y

OpenStudy (anonymous):

so add 5y and subtract 15 from both sides, you get 3x - 15 = 5y, now divide by 5, you get 3/5x - 3 = y

OpenStudy (anonymous):

3 - 3x /y

OpenStudy (anonymous):

read that wrong....follow camp

OpenStudy (anonymous):

-y = -3x/5-3

OpenStudy (anonymous):

5y = 3x-15 y = (3x-15)/5

OpenStudy (anonymous):

see first of all transpose 3x to right hand side -5y = 15 - 3x divide both sides by ( -5) -5y/(-5) = (15-3x)/(-5) y = (15-3x)/(-5) y= (-15+3x)/5

OpenStudy (anonymous):

jimmy you good with calc?

OpenStudy (anonymous):

got it eriklkl

OpenStudy (anonymous):

u can ask more questions if u have

OpenStudy (anonymous):

3x-5y=15 Since 3x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3x from both sides. -5y=-3x+15 Divide each term in the equation by -5. -(5y)/(-5)=-(3x)/(-5)+(15)/(-5) Simplify the left-hand side of the equation by canceling the common terms. y=-(3x)/(-5)+(15)/(-5) Simplify the right-hand side of the equation by simplifying each term. y=(3x)/(5)-3

OpenStudy (anonymous):

it was more easier solution hope u got it eriklkl